Find the average value of the function f(x) = x3 + 4 on the interval [2,4].
Answer #1

f(x)=x^3+4

Interval [2,4]

Average value of the function

f_{avg}=\frac{1}{b-a}\int_{a}^{b}f(x)dx

=\frac{1}{4-2}\int_{2}^{4}(x^3+4)dx

=\frac{1}{2}\int_{2}^{4}(x^3+4)dx

=\frac{1}{2}[\frac{x^{3+1}}{3+1}+4x]_{2}^{4}

=\frac{1}{2}[\frac{x^4}{4}+4x]_{2}^{4}

=\frac{1}{2}[\frac{(4)^4}{4}+4(4)-(\frac{(2)^4}{4}+4(2))]

=\frac{1}{2}[68]

\boldsymbol{=34}